# Graph Reinforcement Learning in Power Grids: A Survey

### v1

@@ Abstract
The challenges posed by renewable energy and distributed electricity generation motivate the development of deep learning approaches to overcome the lack of flexibility of traditional methods in power grids use cases. The application of Graph Neural Networks is particularly promising due to their ability to learn from graph-structured data present in power grids. Combined with Reinforcement Learning (RL), they can serve as control approaches to determine remedial grid actions. This review analyses the ability of Graph Reinforcement Learning (GRL) to capture the inherent graph structure of power grids to improve representation learning and decision making in different power grid use cases. It distinguishes between common problems in transmission and distribution grids and explores the synergy between RL and GNNs. In transmission grids, GRL typically addresses automated grid management and topology control, whereas on the distribution side, GRL concentrates more on voltage regulation. We analyzed the selected papers based on their graph structure and GNN model, the applied RL algorithm, and their overall contributions. Although GRL demonstrate adaptability in the face of unpredictable events and noisy or incomplete data, it primarily serves as a proof of concept at this stage. There are multiple open challenges and limitations that need to be addressed when considering the application of RL to real power grid operation.
@@ 1 Introduction
The role of electrical grid operators, both for transmission and distribution grids, is to ensure cost-efficient availability at all times. However, power systems worldwide are undergoing a paradigm shift driven by the need for CO2 neutrality. The incorporation of renewable distributed generation, and additional load demand due to heating and traffic sector electrification introduce complexities that traditional power system operation struggles to address. These trends necessitate advanced methodologies for optimal operation Marot et al., (2021); Kelly et al., (2020). Although the ongoing energy transition poses major challenges to grid operators, it concerns other stakeholders as well. Markets must adapt to the decentralized structure while new market players emerge, such as Electric Vehicle (EV) charging operators. Further, with the ongoing digitization and build-up of communication systems, the classical power system transforms into a cyber-physical energy system (CPES) Steinbrink et al., (2018). All these new challenges introduce a new layer of complexity into overall power grid operation, which has to be considered for the success of the energy transition.
@@
Traditionally, grid operation has mostly relied on traditional optimization approaches for Optimal Power Flow (OPF) problems. Due to the non-linear and non-convex nature of OPFs, grid operators often fall back on relaxation techniques or heuristic approaches Srivastava et al., (2023). However, relaxation techniques may yield imprecise results, and cannot ensure the solution’s optimality, thus casting doubt on their effectiveness in application. Additionally, classical optimization approaches struggle to scale to real power grids such that exact results can be achieved in a reasonable time, thus relaxation techniques are applied Molzahn and Hiskens, (2019). Noisy or missing measurements cannot reliably be handled by classical approaches Liu et al., (2020). Therefore, practitioners are are exploring deep-learning solutions Marot et al., (2020); Marot et al., 2022a for problems involving power flows such as grid operation. Such solutions appear to be a promising alternative to classical approaches, addressing the challenges of time criticality, scalability and reliability of results.
@@
Deep Reinforcement Learning (DRL) techniques can identify and utilize under-exploited flexibilities in power grids, often overlooked by traditional methods and human operators Kelly et al., (2020). By learning from interactions with the grid environment, DRL agents can dynamically adjust to changing conditions and unforeseen events, potentially preventing cascading failures and blackouts Marot et al., (2021); Donnot et al., (2017). Furthermore, its capacity to consider long time horizons aligns with the inherently dynamic nature of power grids Viebahn et al., 2022a . However, the development and training of DRL agents require extensive simulations, as direct interaction with the physical grid is impractical. These simulations often need to abstract from reality and rarely use real data, leading to challenges in transferring the solutions back to real-world application Kaspar et al., (2020). Furthermore, the challenge of large combinatorial action spaces in power grids hinders the direct application of DRL Dulac-Arnold et al., (2021), highlighting the need for handcrafted action spaces and other reduction techniques EI Innovation Lab, (2020); Lehna et al., (2022, 2024). Despite these challenges, the potential for RL to transform power grid management is significant, particularly as power systems aim at meeting decarbonization goals and accommodating more renewable energy sources Prostejovsky et al., (2019); Marot et al., 2022c . Rather than replacing human operators, the goal is to provide RL-driven action recommendations for themViebahn et al., 2022a ; Marot et al., 2022b . While proof-of-concept results are impressive, research on DRL for power grid control is still in its early stages and a significant gap remains before deployment can be achieved.
@@
The complex interconnectedness within a power grid can naturally be modelled as a graph, in which nodes and edges correspond to elements and connections in the grids Viebahn et al., 2022a . Since GNNs are specifically designed for such graph-structured data, they are highly suitable for modelling interdependencies in power systemsLiao et al., (2021). Furthermore, they excel at capturing spatial relationships and enable an effective feature extraction from the grid. While feed-forward neural networks struggle to produce accurate results when the grid’s topology, and thus the input dimensionality, changes, GNNs are more robust to modifications of the graph structure. This is an advantage, as several grid actions such as bus-bar splitting can transform a single node in a graph into two distinct nodes (or alternatively, combine two nodes into one). This is uncommon in other types of networks Donon et al., (2020). The distinctive configuration of substations requires tailored methods to account for specific network constraints and leverage network flexibilities. Similarly, efficiently managing long-range dependencies such as non-local effects and the rapid propagation speed of electricity pose unique challenges for power systems. In other words, the effectiveness of GNNs across various topologies and different power grids still requires extensive research Ringsquandl et al., 2021a . Furthermore, the interpolation capabilities of GNNs are instrumental in reconstructing missing information and smoothing out noisy measurements Kuppannagari et al., (2021). This addresses practical issues in power systems where sensors may experience connectivity problems, leading to incomplete or unreliable data.
@@
The combination of GNNs and RL represents a synergy that harnesses the strengths of both paradigms. GNNs provide a powerful tool for feature extraction and representation in graph-structured data, enhancing the RL agent’s understanding of the complex relationships within power grids. As pointed out by Munikoti et al., (2023), the performance of RL agents strongly depends on the state encoding, and GNNs are much better encoders for graph-structured environments. Incorporating them into RL has the potential to enhance the agent’s ability to learn from the graph structure. This has the potential for more informed decision-making, better adaptability to changing network conditions, and improved generalization across diverse scenarios and topologies.
@@
This survey explores recent trends in leveraging GRL for power grids, specifically focusing on transmission and distribution networks as well as other power grid applications such as the energy market, communication networks within power systems, and EV charging management. We focus on DRL models that utilize GNNs to capture the graph-structured state space of power grids while leveraging DRL for sequential decision making.
1.1 Adressed Use Cases
1.1.1 Transmission Grid Control
To ensure safe and reliable transmission, human experts traditionally manage and monitor the power grid in control rooms manually. However, the integration of stochastic and dynamic elements, such as renewable energy sources increases the complexity of power grid operation, starting to surpass human expertise Marot et al., (2021). Therefore, automation and data-driven optimization become necessary. The challenge lies in automatically controlling large-scale power grids, which requires intricate yet dependable decision-making in a large decision space.
Many approaches for transmission grid management focus on controlling electricity generation, or loads, e.g., re-dispatch Kamel et al., (2020); Bai et al., (2023); Fuxjäger et al., (2023) or load shedding Larik et al., (2019). An effective alternative solution can be achieved by managing the power grid through topology control, such as substation switching. This grid reconfiguration allows for efficient rerouting of power flow while using existing infrastructure. This is less costly than generator re-dispatch and minimizes disruptions to the existing system. Despite preliminary studies on grid topology control in literature, the large action spaces, and non-linear nature of the problem hinder the practical deployment of existing methods in real-world scenarios.
Researchers have explored multiple approaches to optimize power flow on the popular Grid2op Donnot, (2020) environments of the Learning to Run a Power Network (L2RPN) challenges Marot et al., (2020). Such a simulation and testbed is crucial for developing methods to tackle congestion and enhance grid reliability in networks at different scales. Therefore, research on RL for topology control so far is dominated by Grid2op, highlighting its pivotal role in advancing the field.
1.1.2 Voltage Control for Distribution Grids
The distribution side of power systems, characterized by the need for decentralized decision-making, presents its own challenges. The integration of distributed generation and ramping up the connection of new loads, such as EVs and heat pumps, poses challenges in voltage control for distribution grids Srivastava et al., (2023). Additionally, more sensors also come with an increased amount of available data, which can improve the training of machine learning models such as GRL. In contrast to heuristics, such approaches can use available data from multi-model sources without major effort. Furthermore, the exploration of more comprehensive optimization problems, which use various voltage control options within distribution grids, such as the integration of novel assets or leveraging the functionalities of distributed energy management systems, are promising in the advancement of distribution grid technology. In this context, RL is particularly suitable as it can deal with multiple objectives at the same time.
1.1.3 Other Use cases
GRL shows potential for related applications such as new energy markets, communication networks for power operation, and cross-sectoral applications like EV charging scheduling with power grid constraints. In energy markets, traditional centralized methods struggle with privacy and scalability, making distributed methods like multi-agent RL necessary. GNNs have the potential to enhance these methods by improving modeling accuracy and scalability. In addition to the physical transmission of power, modern energy systems rely on efficient information transmission for monitoring and control. This requires optimized communication networks to route packets effectively without losing critical data. GRL shows promise to optimize these networks, supporting physical grid operations. Finally, GRL is also promising in addressing challenges from the rapid growth of electromobility by determining charging station recommendations and navigation to address power grid boundaries and constraints, thus aiding the integration of electromobility into the broader energy landscape.
@@ 1.2 Contribution and Structure
  • We are the first to comprehensively analyze existing GRL approaches for different power grid use cases, categorizing them based on the specific scenarios they address. Our analysis particularly emphasizes applications in distribution and transmission grids. Within the distribution grid, we differentiate between regular voltage control and emergency situations, while for the transmission grid, we concentrate on topology control and the relevant frameworks.
@@
  • We outline the particular advantages of GNNs for RL in power systems, and identify gaps and open problems in existing approaches.
@@
The papers we analyse in this survey are all published after 2020 up until May 2024, as GRL is a relatively new field. Our selection includes the papers that explicitly address power grids using the combination of GNNs and RL. In particular, we focus on grid operation, which is typically approached for distribution and transmission grids, but we also consider use cases in energy markets, EV applications, and power communication networks if they consider the underlying power grid in their approach.
This review is structured as follows: In Ch. 2, we outline related reviews that deal with methodologies or use cases close to ours. Second, we present the basics of transmission and distribution grids, as well as GNNs and RL in Ch. 3, including the most important concepts used in the analysed approaches. Then, in the main part of this review, we discuss the presented methods in detail and group them by the use case addressed. This part starts with approaches for common problems in transmission grids (Ch. 4) and continues with applications that are typically relevant in distribution grids (Ch. 5). Then, in Ch. 6, we highlight a few relevant papers that cater to different related use cases, such as EV charging. Finally, we give an overall conclusion and outlook.
@@ 2 Related Work
In the field of RL as well as GNNs, several more general surveys have been published highlighting different aspects of the respective method or focusing on specific application areas. Broader GNNs surveys that cover the most common GNN architectures are for example presented by Zhou et al., (2020), Thomas et al., (2022) or Wu et al., (2020). Skarding et al., (2021) focus on GNNs for dynamic graphs that evolve over time, while Wu et al., 2022b is an application specific survey that examines GNN approaches for recommender systems. Similarly, in the field of RL, there are several general reviews such as Arulkumaran et al., (2017); Garcıa and Fernández, (2015); Zhu et al., (2023).
Regarding DRL in the context of energy systems, Zhang et al Zhang et al., (2019) provide an overview of existing approaches. They analyze different DRL algorithms that address energy-related problems such as energy management, demand response, electricity market or operational control. There are also more specialized works, such as the review by Vázquez-Canteli et al. Vázquez-Canteli and Nagy, (2019) that examines approaches for RL approaches that address demand response applications in smart grids. However, these works do not include graph reinforcement learning, since it is not yet as well established as RL combined with traditional neural networks and the number of publications has only increased in recent years.
Surveys or literature reviews of approaches combing these two methods are rather rare also in other application areas. Munikoti et al. Munikoti et al., (2023) present a general survey on GRL identifying 80 relevant papers on different use cases. They divide the approaches into two categories: DRL enhancing GNNs and GNNs enhancing DRL. The former deals with DRL applied to architecture search, improving the explainability of GNNs or generating adversarial attacks on GNNs. The latter one focuses on DRL approaches that use GNNs within a DRL setup, which is also the focus of this review. The application areas they investigate are combinatorial optimization, transportation, control of system such wireless networks or material flow.
Fathinezhad et al., (2023) present another survey on GRL that focuses on the basics of GNNs and RL and their relationship. It points out the concepts of how graphs and RL are combined and does not investigate any specific application areas. Furthermore, the authors particularly focus on multi-agent RL, where GNNs are used to enable communication and information exchange between agents. We, in contrast, focus more on the use of GNNs as feature extractors for data that has an inherent graph structure. Fathinezhad et al., (2023) mention only one energy-related application (Pei et al., (2023)), but do not analyze it in detail, since application-specific approaches are not the focus of their survey.
The survey presented by Nie et al., (2023), also examines GRL from a more methodological point of view. They analyze in detail, how RL can be used to improve GNNs and to solve graph problems in general. The survey covers several application fields which are transportation network optimization, E-commerce recommendation systems, structure prediction and generation in medical research, and modelling and controlling of the COVID19-virus.
The survey presented by Mazyavkina et al., (2021) explores GNNs as representation learners for graph structures problem in the RL setup, however the authors concentrate exclusively on on combinatorial optimization problems.
In Mendonca et al., (2019) give a detailed overview of how different graph algorithms can be used to enhance RL by abstracting primitive actions into temporal macro-actions. Similarly, Pateria et al., (2021) analyze hierarchical RL which often involves graph-based approaches, e.g. to discover subtasks, but both works do not consider GNNs in that context.
In all of the above mentioned related works, the application to the energy system is not reviewed. A review focusing on GNNs and their application in power systems is presented by Liao et al., (2021). This review, published in 2022, analyses different application-specific approaches to fault scenarios, time series prediction for renewable energy, data generation and power flow calculation with respect to the GNN architecture used. Liao et al. point out that the reviewed GNN methods typically show better performance than traditional neural networks in the experiments conducted. However, there are still several gaps and open questions that require further investigation, especially due to the novelty of the field. Liao et al. do not consider approaches using RL.
Regarding the different use cases in power grids, there are related works that specialise in specific problems such as voltage control, e.g. Srivastava et al., (2023) or Murray et al., (2021). However, the works considered are mainly traditional optimisation methods and heuristics, since such approaches are still the most dominant in practice. To the best of our knowledge, a detailed overview on the application of GRL in the context of energy systems do not yet exist. Therefore, we aim to fill this gap with this work.
@@ 3.1 Power Grids
Power grids are essential to modern society and a crucial part of today’s infrastructure. In the face of the energy transition, power system engineers encounter challenges across all aspects of the grid. Its general purpose is to transport electricity from generation units to consumers, which typically are not located in the same area. Traditionally, generation was centralized, e.g., at fossil-fueled or nuclear power plants, resulting in unidirectional power flow from generation to consumption. With the worldwide rise of renewables, generation units are spread across the power system resulting in a decentralized structure and bi-directional power flow.
With the electrification of the traffic and heating sectors, the consumption side of power grids is also undergoing a major shift. Not only does the overall consumption increase with this sector coupling, but also new simultaneity of consumption have to be considered by power system engineers, e.g., simultaneous heating in winter or swing season resulting in a simultaneous heat pump operation over a grid area.
Power grids are typically divided into two levels: transmission and distribution. These levels differ in voltage levels, purpose, and the resulting characteristics, and are split by substations (see 1). We elaborate on both power grid levels in the following subsections.

### v3

# Graph Reinforcement Learning for Power Grids: A Comprehensive Survey

@@ Abstract
The rise of renewable energy and distributed generation requires new approaches to overcome the limitations of traditional methods. In this context, Graph Neural Networks are promising due to their ability to learn from graph-structured data. Combined with Reinforcement Learning, they can serve as control approaches to determine remedial network actions. This review analyses how Graph Reinforcement Learning (GRL) can improve representation learning and decision making in power grid use cases. Although GRL has demonstrated adaptability to unpredictable events and noisy data, it is primarily at a proof-of-concept stage. We highlight open challenges and limitations with respect to real-world applications.
@@ 1 Introduction
The role of electrical grid operators, both for transmission and distribution grids, is to ensure cost-efficient availability at all times. However, power systems worldwide are undergoing a paradigm shift driven by the need for CO2 neutrality. The integration of renewable distributed generation and additional load demand due to heating and traffic sector electrification introduce complexities that traditional power system operations are struggling to cope with. These trends require advanced methods for optimal operation Marot et al., (2021); Kelly et al., (2020). The ongoing energy transition also affects other stakeholders, such as energy market participants. They need to adapt to the decentralized structure and new market players such as Electric Vehicle (EV) charging operators. Additionally, the ongoing digitization and build-up of communication systems transform the classical power system into a cyber-physical energy system (CPES) Steinbrink et al., (2018). All these new challenges introduce a new layer of complexity to power grid operation.
@@
Traditionally, grid operation has mostly relied on optimization approaches for Optimal Power Flow (OPF) problems. However, due to the non-linear and non-convex nature of OPFs, these approaches struggle to scale to real power grids such that exact results cannot be achieved in a reasonable time Srivastava et al., (2023). Therefore, relaxation techniques are used to reduce the complexity Molzahn and Hiskens, (2019). However, they can produce imprecise results and cannot ensure optimality, which casts doubt on their effectiveness in application.
Furthermore, noisy or missing measurements cannot be reliably handled by classical approaches Liu et al., (2020). Therefore, practitioners are exploring deep learning solutions Marot et al., (2020); Marot et al., 2022a for power flow problems. Such solutions are promising alternatives to classical approaches, addressing the challenges of time criticality, scalability, and reliability of results.
@@
Deep Reinforcement Learning (DRL) techniques can identify and exploit under-utilized flexibilities in power grids, often overlooked by traditional methods and human operators Kelly et al., (2020). By learning from interactions with the grid environment, DRL agents can dynamically adjust to changing conditions and unforeseen events, potentially preventing cascading failures and blackouts Marot et al., (2021); Donnot et al., (2017). Furthermore, their ability to consider long time horizons aligns with the dynamic nature of power grids Viebahn et al., 2022a . However, the development and training of DRL agents requires extensive simulations, as direct interaction with the physical grid is impractical. These simulations often need to abstract from reality and rarely use real data, leading to challenges in transferring the solutions back to real-world applications Kaspar et al., (2020). Furthermore, the large combinatorial action spaces in power grids hinder the application of DRL Dulac-Arnold et al., (2021), highlighting the need for handcrafted action spaces and other reduction techniques EI Innovation Lab, (2020); Lehna et al., (2022); Lehna et al., 2024b . Despite these challenges, the potential for Reinforcement Learning (RL) in power grid management is significant, particularly as power systems aim at meeting decarbonization goals Prostejovsky et al., (2019); Marot et al., 2022c . The aim is not to replace human operators but to provide them with RL-driven action recommendations Viebahn et al., 2022a ; Marot et al., 2022b . Despite impressive proof-of-concept results, DRL research for grid control is still at an early stage and there are significant gaps to be filled before deployment.
@@
Power grids can naturally be modeled as a graph, in which nodes and edges correspond to grid elements and their connections Viebahn et al., 2022a . Since Graph Neural Networks are specifically designed for such graph-structured data, they are highly suitable for modeling interdependencies in power systems Liao et al., (2021). Furthermore, they can capture relationships between elements and enable an effective feature extraction from the grid. While feed-forward neural networks struggle to produce accurate results when the grid’s topology, and thus the input dimensionality, changes, GNNs are more robust to modifications of the graph structure. This is an advantage, as several grid actions, such as bus-bar splitting, can transform a single node in a graph into two distinct nodes (or alternatively, combine two nodes into one). This is uncommon in other types of networks Donon et al., (2020) and thus requires tailored methods. Similarly, long-range dependencies such as non-local effects and the rapid propagation speed of electricity pose unique challenges. Therefore, the effectiveness of GNNs across various topologies and different power grids still require extensive research Ringsquandl et al., 2021a . Furthermore, the interpolation capabilities of GNNs are instrumental in reconstructing missing information and smoothing out noisy measurements Kuppannagari et al., (2021). This addresses practical issues in power systems where sensors may experience connectivity problems, leading to incomplete or unreliable data.
@@
The combination of GNNs and RL represents a synergy that harnesses the strengths of both paradigms. GNNs provide a powerful tool for feature extraction in graph-structured data, enhancing the RL agent’s understanding of the complex relationships within power grids. As pointed out by Munikoti et al., (2023), the performance of RL agents strongly depends on the state encoding, and GNNs are much better encoders for graph-structured environments. Incorporating them into RL has the potential for more informed decision-making, better adaptability to changing network conditions, and improved generalization across diverse scenarios and topologies.
@@
This survey explores recent trends in Graph Reinforcement Learning (GRL) for power grids, specifically focusing on transmission and distribution grids as well as other power grid applications such as the energy market, communication networks within power systems, and EV charging management. We focus on DRL models that utilize GNNs to capture the graph-structured state space of power grids while leveraging DRL for sequential decision-making.
@@ 1.1 Contribution and Structure
  • We are the first to comprehensively analyze existing GRL approaches for different power grid use cases, categorizing them based on the specific scenarios they address. Our analysis emphasizes applications in distribution and transmission grids. Within the distribution grid, we differentiate between regular voltage control and emergency situations, while for the transmission grid, we concentrate on topology control and the relevant frameworks.
@@
  • We outline the particular advantages of GNNs for RL in power systems and identify gaps and open problems in existing approaches.
@@
The papers we analyze are all published between 2020 and May 2024, as GRL is a relatively new field. Our selection includes those papers that explicitly address power grids using the combination of GNNs and RL. In particular, we focus on grid operation in distribution and transmission grids, but we also consider use cases in energy markets, EV applications, and communication networks if they consider the underlying power grid in their approach.
This review is structured as follows: In Ch. 2, we outline reviews that address methodologies or use cases close to ours. Second, we present the basics of transmission and distribution grids, GNNs and RL in Ch. 3. In the main part of this review, we discuss the presented methods in detail and group them by the use case addressed. This part starts with approaches for common problems in transmission grids (Ch. 4) and continues with applications in distribution grids (Ch. 5). Then, in Ch. 6, we highlight a few relevant papers that cater to different related use cases, such as EV charging. Finally, we give an overall conclusion and outlook.
@@ 2 Related Work
Several general surveys on GNNs cover various methodologies and applications. For example, Zhou et al., (2020), Thomas et al., (2022), and Wu et al., (2020) provide overviews of common architectures, while Skarding et al., (2021) focus on dynamic graphs that evolve over time, and Wu et al., 2022b examines GNNs in recommender systems. As for the application area of power systems, Liao et al., (2021) reviews GNNs, highlighting their superior performance over traditional neural networks but noting gaps and open questions, particularly as RL is not covered.
Similarly in RL, general reviews include Arulkumaran et al., (2017), Garcıa and Fernández, (2015), and Zhu et al., (2023). Zhang et al., (2019) explores DRL for energy systems, focusing on problems like demand response and electricity market, or operational control. Specialized reviews, such as Vázquez-Canteli and Nagy, (2019), focus on demand response in smart grids but do not cover GNNs, which is a more recent development compared to traditional RL approaches.
Literature reviews combining GNN and RL are rare. Munikoti et al., (2023) survey 80 relevant papers, categorizing them into DRL-enhancing GNNs and GNNs-enhancing DRL. The former includes DRL for architecture search and improving GNN explainability, while the latter covers GNN use in DRL, which is closer to our work. They explore areas like combinatorial optimization and transportation but exclude energy applications.
Fathinezhad et al., (2023) survey GRL approaches with a focus on the methodology of GNNs and RL, especially in multi-agent settings where GNNs facilitate agent communication. They primarily explore how graphs and RL interact, while our focus is on using GNNs as feature extractors for graph-structured power grid data. Although they briefly mention an energy-related application Pei et al., (2023), it is not analyzed in detail as their review does not emphasize application-specific approaches.
Similarly, the survey presented by Nie et al., (2023) examines GRL methodologically, detailing how RL can enhance GNNs and address graph problems. They cover various transportation and medical research applications but do not address energy-related use cases.
Several surveys focus on specific aspects of GRL: Mazyavkina et al., (2021) examine GNNs for representation learning in combinatorial optimization within RL. Mendonca et al., (2019) detail how graph algorithms can enhance RL through action abstraction, while Pateria et al., (2021) explore hierarchical RL with graph-based approaches for discovering subtasks. None of these surveys address power grids.
Related works analyze specific power grid problems such as voltage control, such as Srivastava et al., (2023) and Murray et al., (2021), but they mostly use traditional optimization and heuristics. There is currently no comprehensive overview of GRL applications in power grids. Therefore, we aim to fill this gap with this work.
@@ 3.1 Power Grids
Power grids are essential to modern society and a crucial part of today’s infrastructure. In the face of the energy transition, power system engineers encounter challenges in all aspects of the grid. Their purpose is to transport electricity from generation units to consumers, which are typically not located in the same area. Traditionally, generation has been centralized, e.g., at fossil-fueled or nuclear power plants, resulting in unidirectional power flow from generation to consumption. With the worldwide expansion of renewables, generation units are spread across the power system resulting in a decentralized structure and a bi-directional power flow. With the electrification of the traffic and heating sector, the consumption side of power grids is also undergoing a major shift.
Power grids are typically divided into two levels: transmission and distribution. These levels are split by substations (see Fig. 1) and differ in voltage levels, purpose, and characteristics. We elaborate on both levels in the following subsections.